OpenStudy (anonymous):

What is the exact value of the square root of the quantity 27 over 64.? 3 the square root of 3 all over 4. 9 the square root of 3 all over 4. 9 the square root of 3 all over 8. 3 the square root of 3 all over 8.

OpenStudy (anonymous):

\[\sqrt{\frac{27}{64}}=\frac{\sqrt{27}}{\sqrt{64}}\] is a start the denominator is easy, right?

OpenStudy (anonymous):

i.e. what is the square root of 64?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

ok good so you can ignore answers 1 and 2 denominator is 8

OpenStudy (anonymous):

as for \(\sqrt{27}\) , note that \(27=9\times 3\) and so \(\sqrt{27}=\sqrt{9\times 3}=\sqrt{9}\sqrt{3}\) what is \(\sqrt{9}\) ?

OpenStudy (anonymous):

3 so D

OpenStudy (cggurumanjunath):

3 the square root of 3 all over 8.

OpenStudy (anonymous):

yes, D